OpenStudy (anonymous):

differentiate the following y(t0=(1)/(t+4e^t)....i differentiate f'(x)=1 g'(x)=1+4e^t?

OpenStudy (anonymous):

hold the phone

OpenStudy (anonymous):

is this \[y=\frac{1}{1+4e^t}\]?

OpenStudy (anonymous):

is so no quotient rule needed. the derivative of \[\frac{1}{f(x)}=\frac{-f'(x)}{f^2(x)}\]

OpenStudy (anonymous):

should get \[\frac{-1-4e^t}{(t+4e^t)^2}\] in an instant

OpenStudy (anonymous):

im sorry i went away to class. but thanks i appreciate it